# Tensorized Ant Colony Optimization for GPU Acceleration

### 07 Apr 2024 v1

@@ Abstract.
copyright: acmlicenseddoi: 10.1145/nnnnnnn.nnnnnnnisbn: 978-x-xxxx-xxxx-x/YY/MMconference: The Genetic and Evolutionary Computation Conference 2024; July 14–18, 2024; Melbourne, Australiajournalyear: 2024ccs: Computing methodologies Vector / streaming algorithmsccs: Mathematics of computing Combinatorial optimization
@@ 1. Introduction
The JAX framework (Bradbury et al., 2018) enhances high-performance numerical computing with hardware acceleration significantly. Building upon JAX, pioneering platforms like EvoJAX (Tang et al., 2022), evosax (Lange, 2023), and EvoX (Huang et al., 2023) have emerged, providing efficient tools for evolutionary computation for advanced hardware acceleration. However, a complete implementation of scalable ACO remains absent. To bridge this gap, we propose the Tensorized Ant Colony Optimization (TensorACO) for GPU Acceleration.
@@
  • We propose an Adaptive Independent Roulette (AdaIR) method which improves performance while retaining the benefits of parallelization by dynamically adjusting the tendency of ants to select cities with higher probabilities.
@@ 2. Method
The schematic overview of TensorACO is shown in Figure 1. TensorACO encompasses several key components: Preprocessing, Ant System Tensorization, Ant Path Tensorization, and an Adaptive Selection Mechanism. The symbol $n$ represents the number of cities in TSP, and $m$ represents the number of ants in ACO.
@@ 3. Experiments
All experiments were conducted on a uniform platform and compatible with the EvoX (Huang et al., 2023) framework. (CPU: Intel Core i7-10750H @ 2.60GHz; GPU: NVIDIA A100 Tensor Core; Python 3.9; CUDA 12.2; JAX 0.4.23). The dataset employed in this experiment is the TSPLIB dataset (Reinelt, 1991).

### 12 Apr 2024 v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024copyright: rightsretainedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2024; Melbourne, VIC, Australiabooktitle: Genetic and Evolutionary Computation Conference (GECCO ’24 Companion), July 14–18, 2024, Melbourne, VIC, Australiadoi: 10.1145/3638530.3654394isbn: 979-8-4007-0495-6/24/07ccs: Computing methodologies Vector / streaming algorithmsccs: Computing methodologies Discrete space searchccs: Mathematics of computing Combinatorial optimization
@@ 1. Introduction
The JAX framework (Bradbury et al., 2018) enhances high-performance numerical computing with hardware acceleration significantly. Building upon JAX, pioneering platforms like EvoJAX (Tang et al., 2022), evosax (Lange, 2023), and EvoX (Huang et al., 2024) have emerged, providing efficient tools for evolutionary computation for advanced hardware acceleration. However, a complete implementation of scalable ACO remains absent. To bridge this gap, we propose the Tensorized Ant Colony Optimization (TensorACO) for GPU Acceleration.
@@
  • We propose an Adaptive Independent Roulette (AdaIR) selection mechanism which improves performance while retaining the benefits of parallelization by dynamically adjusting the tendency of ants to select cities with higher probabilities.
@@ 2. Method
The schematic overview of TensorACO is shown in Figure 1. TensorACO encompasses several key components: Preprocessing, Ant System Tensorization, Ant Path Tensorization, and an Adaptive Independent Roulette selection mechanism. The symbol $n$ represents the number of cities in TSP, and $m$ represents the number of ants in ACO.
@@ 3. Experiments
All experiments were conducted on a uniform platform and compatible with the EvoX (Huang et al., 2024) framework. (CPU: Intel Core i7-10750H @ 2.60GHz; GPU: NVIDIA A100 Tensor Core; Python 3.9; CUDA 12.2; JAX 0.4.23). The dataset employed in this experiment is the TSPLIB dataset (Reinelt, 1991).